The value of a piece of equipment has a decay factor of 0.080 per year. After 5 years, the equipment is worth $98,304. What was the original value of the equipment?
Answer by Alan3354(69443)

The value of a piece of equipment has a decay factor of 0.080 per year. After 5 years, the equipment is worth $98,304. What was the original value of the equipment?
------------
Its value each year is 0.92 of the previous year.
After 5 years,
Value = OV*0.92^5 = 98304
OV = 98304/(0.92^5)
OV =~ $149153.02